Remington Corporation is preparing its cash budget for June. The budgeted beginning cash balance is $28,000. Budgeted cash receipts total $140,000. Cash disbursements include raw materials purchases of $35,000, property insurance of $14,500, utilities of $10,000, executive salaries of $79,000, and office supplies of $6,300. The desired ending cash balance is $20,000. The excess (deficiency) of cash available over disbursements for June is: Question 16 options: $23,200 $17,900 $64,000 $21,000
Answer #1

Answer

  • Correct Answer = Option #1: $ 23,200
  • Working

A

Beginning Cash balance

$28,000

B

Cash receipts

$140,000

C = A+B

Total Cash available

$168,000

D = 35000 + 14500 + 10000 + 79000 + 6300

Cash Disbursements

$144,800

E = C - D

Excess of cash available over disbursement

$23,200 Answer

F

Minimum required balance

$20,000

G = F - E

Amount to be borrowed

$ 0